OpenStudy (wilkes688):

find the gcf of the follow monomials 48b^2 21b

OpenStudy (welshfella):

first deal with the 2 numbers 21 and 48 can you think of the greatest number which will divide into these?

OpenStudy (welshfella):

one way to do this is to take out duplicates of their factors 21 = 3*7 48 = 2*2*2*2*3

OpenStudy (welshfella):

what number or numbers is common to both factors?

OpenStudy (welshfella):

- that is your GCF